OpenStudy (anonymous):

a piece of land in the shape of parallelogram whose altitude is one third its base if the perimeter is 200 m and its area is numerically equal to 8 times its perimeter find the length of diagonals

OpenStudy (anonymous):

why base is 3x?

OpenStudy (nurali):

altitude is one third its base

OpenStudy (anonymous):

base should be x and the altitude 1/3x?

OpenStudy (nurali):

altitude = x then base = x/3, area = x/3 * x x^2 = 8*200*3 x^2=4800 x=40sqrt(3)

OpenStudy (anonymous):

ok so I hav now the area, how can I found the diagonals?

OpenStudy (nurali):

d^2=sqrt(200^2+(40sqrt(3)^2) d^2=sqrt(40000+4800) d=80sqrt(7)